OpenStudy (anonymous):

what is the volume of a cylinder with a diameter of 10cm and depth 12cm?

OpenStudy (raden):

Volume cylinder = pi * r^2 ^ height given diameter = 10 then the radius = 1/2 * 10 = 5 cm its height = 12 cm therefore, V = 3.14 * 5^2 * 12 V = 3.14 * 25 * 12 V = 942 cm^3